OpenStudy (anonymous):

In Equation such as 6x-y=29 and then i have x=y-9 i sub in y-9 6(y-9)-y=-29 solve for y which then equals 6y-54-y=-29 5y=25 ( how does it get their from 6y-54-y=-29?)

OpenStudy (anonymous):

how can you get 5y=25 from that and why. i need to find the value of y but i dont understand after the point of 6y-54-y=-29 that you then get 5y=25

OpenStudy (anonymous):

is it 29 or -29 in the 1st eq?

OpenStudy (anonymous):

-29

OpenStudy (anonymous):

Well you have 6y - 54 - y on the left side. And you can rewrite that by the commutative property to be 6y -y - 54 Then you can combine the first two terms 6y-y = 5y so you get 5y - 54 and that equals -29

OpenStudy (anonymous):

With me so far?

OpenStudy (anonymous):

yes so i would subtract 1y to each side?

OpenStudy (anonymous):

No, you add 54 to both sides.

OpenStudy (anonymous):

like 6y- 1y = 5y-54

OpenStudy (anonymous):

Wait what?

OpenStudy (anonymous):

I think you missed something

OpenStudy (anonymous):

You have this: \[\text{ ( given ) }6y - 54 - y = -29\]\[\text{(commute)}\implies 6y - y-54 = -29\]\[\text{(combine first two terms)}\implies 5y - 54 = -29\]

OpenStudy (anonymous):

ohhh i get it

OpenStudy (anonymous):

Because 6 y's minus 1 y is 5 y's. Just like 6 oranges minus 1 orange is 5 oranges.

OpenStudy (anonymous):

okay. so anything other equation like that i should combine like terms

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thank you